(a) A highway patrolman traveling at the speed limit is passed
by a car going 15 mph faster than the speed limit. After one
minute, the patrolman speeds up to 100 mph. How long after
speeding up until the patrolman catches up with the speeding
car. The speed limit is 55 mph.
(b) Same question, but this time the patrolman speeds up to a
speed of v mph (v>70).

(a)_______ minutes
(b)________ minutes

Answer by mananth(12270)

Car speed 70 mph
Patrolman speed 100 mph
Car speed 10:00
Patrolman speed 10:30
Car speed will have covered 1.17 miles before Patrolman speeds up
catch up distance= 1.17 miles
catch up speed = 100 -70 mph
catch up speed = 30 mph
Catchup time = catchup distance/catch up speed
catch up time= 1.17 / 30
catch up time= 0.04 hours 2.33 minutes
catchup speed = v-70
catchup distance = 1.17 miles
time taken = 1.17/(v-70) *60 minutes